OpenStudy (anonymous):

What is the equation of the line that passes through (1, 3) and (-2, -3)? y = -2x + 1 y = 2x + 1 correct? y = 1/2 x - 1 y = -1/2 x + 1

OpenStudy (anonymous):

Find the slope and then plug that into standard form of y = mx+b

OpenStudy (saszombieassaulter):

Slope formula: \(\Large m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\) And of you prefer use the point-slope formula: \(\Large y-y_{1}=m(x-x_{1})\)

OpenStudy (anonymous):

m=y^2-y^1 = -6/-3

OpenStudy (anonymous):

im a little confused

OpenStudy (anonymous):

ok